Evaluate the surface integral

  
S

F · dS

for the given vector field F and the oriented surface S. In other words, find the flux of F across S. For closed surfaces, use the positive (outward) orientation.

F(x, y, z) = x i − z j + y k


S is the part of the sphere

x2 + y2 + z2 = 36

in the first octant, with orientation toward the origin
Answer #1

Given that F = xi-zj+yX divF = *** () + (-2) + 2()=1+0+0=1 Given surface Sis the sphere r* +*+z= 6in the first octant By Gaus